• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name G.Kumbartzki

    Log entry time 16:03:04 on June22,2002

    Entry number 86602

    keyword=Shift Summary

    8:17 Problems with the liq.He system brought all magnets down.

    8:40 LH flow is back, Cryo-levels recovered. Reset of power supplies leaves Quench fault on
    9:28 asked for access to the Hall.
    10:39 Start cycling Q2 and Q3 on left and right specs.
    Since it takes a long time to set the right arm dipole we skip the efficiency run and switch to the next kinematic setting Q2_i40.
    11:26 ready for beam
    Run 1724 11:30 raster off, new trigger set and preset scaling of triggers
    Run 1725 12:10 start production run. For about 40 min had very unstable beam while MCC tried to make max beam. settled around 75 uA. Still we have frequent beam trips (talked to MCC they expierence RF trips in addition to BCM trips)

    13:29 Restart Coda. Beam out, MCC will try to find out what is wrong with the beam

    14:10 Beam is back going up to 95 uA and stays stable.
    Run 1728,1729
    15:12 Coda Restart (twice)
    Left Arm Q2 shut down? Reset, Recycled
    Then L Q1 He lvl Fault#1 and 2 came on and Q1 went down!
    It took the end of the shift to have all up and running again.